[The following information applies to the questions displayed below.]

Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labour-hours, and its standard costs per unit are as follows:

  Direct materials: 5 kg at $10.00 per kg $ 50.00  
  Direct labour: 3 hours at $17 per hour 51.00  
  Variable overhead: 3 hours at $7 per hour 21.00  
  Total standard cost per unit $ 122.00  

The company planned to produce and sell 24,000 units in March. However, during March the company actually produced and sold 30,600 units and incurred the following costs:

  1. Purchased 170,000 kg of raw materials at a cost of $9.00 per kg. All of this material was used in production.
  2. Direct labour: 68,000 hours at a rate of $18 per hour.
  3. Total variable manufacturing overhead for the month was $512,040.

5. What is the labour rate variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Do not round intermediate calculations.)
Answer #1

Labor rate variance

= (SR-AR) *ALH

= (17 - 18)*68,000

= - 1*68,000

= 68,000 Unfavorable